# Novel very-high-frequency quasi-periodic oscillations of compact, non-singular objects

### v1

@@ I Introduction
which we will refer to as the “horizon condition” [18]. If the horizon condition is violated, we refer to such a geometry as horizonless, and the resulting non-singular metric ceases to describe a black hole and describes an exotic, compact object instead.
@@
Notably, such horizonless feature two stable circular orbits: the well-known innermost stable circular orbit (ISCO) at $r=r_{\text{ISCO}}$ that is slightly affected by the presence of the regulator $L$, as well as an additional stable, circular orbit at
@@
which exists for all $L$, but is only part of the exterior spacetime in the absence of the horizon. For compact objects in the stellar mass range the resulting QPO’s from that orbit exceed those from the ISCO position by a factor of $\mathcal{O}(20)$, justifying the name “very-high frequency” QPO’s (VHFQPO’s). In a slight abuse of nomenclature, we still refer to the (formerly innermost) stable circular orbit as the ISCO, whereas we will refer to the additional orbit as L-ISCO (as in $L$-induced innermost stable circular orbit). For a sketch of the shape of such an effective potential and the position of these two orbits see Fig. 1.
Figure 1: We sketch the effective potential of a massive particle in the Schwarzschild geometry (dotted line) and in the spacetime of a non-singular compact horizonless object (solid line). The Schwarzschild geometry features an innermost stable circular orbit (ISCO) at $r=r_{\text{ISCO}}$, and in presence of a regulator $L>0$ its location is shifted slightly. More importantly, in case of the horizonless geometry, a new entirely $L$-induced innermost stable circular orbit (L-ISCO) at $r=r_{\text{L-ISCO}}$ arises. The existence of this L-ISCO is guaranteed by the regularity of the metric of the compact non-singular object, inducing an inflection point in the effective potential at small distances $r\sim L$. Notably, the L-ISCO is only connected to the entire spacetime if $L>L_{\star}$ such that the resulting metric has no horizon and describes a compact, non-singular object.
@@ III.1 Prototypical examples
Table 1: We list the four non-singular black hole metric functions under consideration in this paper, where the metric is given by $\mbox{d}s^{2}=-f(r)\mbox{d}t^{2}+\mbox{d}r^{2}/f(r)+R^{2}(r)\mbox{d}\Omega^{2}$, and we defined the mass function $M(r)$ via the expression f(r)=12GM(r)/rf(r)=1-2GM(r)/r. For each metric, we list the conditions on the regulator for the existence of a black hole horizon.
@@ V Conclusions
There are several ways in which the presented studies can be extended. First, what is the role of angular momentum? Does its presence shift the values of the VHFQPO’s to smaller or larger values? And, second, a detailed analysis of the sensitivity of experimental missions to resolve the existence of potential VHFQPO’s is desirable, along with a more streamlined analysis toolchain involving a Markov chain Monte Carlo analyses as presented e.g. in Ref. [11]. Last, we also recently explored the notion of mass-dependent regulators [18], and it would be interesting to track the impact of such non-standard regulators on the phenomenology of the described VHFQPO’s.

### v2

@@ I Introduction
which we will refer to as the “horizon condition.” If the horizon condition is violated, we refer to such a geometry as horizonless, and the resulting non-singular metric ceases to describe a black hole and describes an exotic, compact object instead [23, 24, 25].
@@
For stationary, non-singular geometries, marginally stable massless orbits (“light rings”) occur in pairs [26, 27]. Similar existence theorems have been formulated for massive orbits [28, 29] that are the subject of the present work. Such horizonless geometries can hence feature two stable circular orbits: the well-known innermost stable circular orbit (ISCO) at $r=r_{\text{ISCO}}$ that is slightly affected by the presence of the regulator $L$, as well as an additional stable, circular orbit at
@@
which exists for all $L$, but is only part of the exterior spacetime in the absence of the horizon. To date, its astrophysical relevance in connection to QPO’s has not been discussed. For compact objects in the stellar mass range the resulting QPO’s from that orbit exceed those from the ISCO position by a factor of $\mathcal{O}(20)$, justifying the name “very-high frequency” QPO’s (VHFQPO’s). In a slight abuse of nomenclature, we still refer to the (formerly innermost) stable circular orbit as the ISCO, whereas we will refer to the additional orbit as L-ISCO (as in $L$-induced innermost stable circular orbit). For a sketch of the shape of such an effective potential and the position of these two orbits see Fig. 1. The non-singular geometry close to the coordinate origin $r=0$ endows the effective potential with an additional minimum at small radii. The slope at this minimum is dictated by the regulator scale $L$ and hence very sensitive to its precise value, making quasi-periodic oscillations around that minimum a potential probe for such new physics. [^2] Inner stable massless orbits raise stability issues due to the trapping of electromagnetic and gravitational radiation [30]. This has since been shown to not generically lead to instabilities but rather introduce a weak high-frequency radiation “hair” [31]. While it is possible that similar conclusion holds for massive orbits, the test particle analysis performed in the present work is not adapted to answer this question. We will therefore comment on the limitations of this analysis in the Conclusions.
Figure 1: We sketch the effective potential of a massive particle in the Schwarzschild geometry (dotted line) and in the spacetime of a non-singular compact horizonless object (solid line) as a function of distance in arbitrary units. The Schwarzschild geometry features an innermost stable circular orbit (ISCO) at $r=r_{\text{ISCO}}$, and in presence of a regulator $L>0$ its location is shifted slightly outward. More importantly, in case of the horizonless geometry, a new entirely $L$-induced innermost stable circular orbit (L-ISCO) at $r=r_{\text{L-ISCO}}$ arises. The existence of this L-ISCO is guaranteed by the regularity of the metric of the compact non-singular object, inducing an inflection point in the effective potential at small distances $r\sim L$. Notably, the L-ISCO is only connected to the entire spacetime if $L>L_{\star}$ such that the resulting metric has no horizon and describes a compact, non-singular object.
@@ III.1 Prototypical examples
Table 1: The four non-singular black hole metric functions under consideration in this paper, where the metric is given by $\mbox{d}s^{2}=-f(r)\mbox{d}t^{2}+\mbox{d}r^{2}/f(r)+R^{2}(r)\mbox{d}\Omega^{2}$. We defined the mass function $M(r)$ via the expression f(r)=12GM(r)/rf(r)=1-2GM(r)/r, and rescale $L_{\text{H}}=\sqrt{L^{3}/(2GM)}$ for the Hayward metric in what follows. For each metric, the last column gives the maximum regulator value $L_{\star}$ for which a black hole horizon exists.
@@ V Conclusions
However, there are limitations to the test particle approach, which we would like to address now briefly.
In this work, the effects of the backreaction of accumulating particles at the L-ISCO (the new, innermost stable circular orbit) onto the geometry have been ignored. It is possible that friction between these particles leads to a natural depletion of angular momentum of massive particles on the L-ISCO, inducing a de-population of the L-ISCO that could prevent potential instabilities from occurring. At the same time, it is also possible that at distances $r\sim L$, where the L-ISCO is located, effects beyond general relativity become important that would lead to a different coupling between matter and gravity or matter among itself.
For a deeper understanding of the L-ISCO VHFQPO phenomenology for compact, horizonless objects it is hence necessary to develop a framework that can account for the backreaction of the background geometry. One possible consequence could be a dampening of the VHFQPO signal, but it is also conceivable that unstable horizonless configurations re-collapse into black holes and disconnect the L-ISCO region from the external spacetime, removing the signals completely from the spectrum. Last, the time scale of such effects is important as well, for if the recollapse happens slowly it is conceivable that a part of the oscillations would escape to infinity and could then be theoretically observable in a reasonably parametrized short-time Fourier transform of the X-ray spectrum.
For these reasons, the reported results on VHFQPO’s should be seen as an upper bound of a possible observational effect that should be refined in future analyses.
Finally, there are several ways in which the presented studies can be extended. First, what is the role of angular momentum? Does its presence shift the values of the VHFQPO’s to smaller or larger values? And, second, a detailed analysis of the sensitivity of experimental missions to resolve the existence of potential VHFQPO’s is desirable, along with a more streamlined analysis toolchain involving a Markov chain Monte Carlo analyses as presented e.g. in Ref. [11]. Last, we also recently explored the notion of mass-dependent regulators [25], and it would be interesting to track the impact of such non-standard regulators on the phenomenology of the described VHFQPO’s.